Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 141     Entry time: 18 Apr 2023, 23:01
Author: Laurent Mercadier, Martin Teichmann 
Group: SCS 
Subject: summary late shift day 1 

Achievements

  • knife edge scans of FEL beam, 30 um x 30 um beam size
  • aligned beam stop in front of camera
  • triggering pattern tested and works, still need to test with laser and XGM

Problems

  • Y motor of beam stop doesn't reach the desired up position
    • had to move the entire camera instead
  • Once or twice scanner Y got into error during knife edge scan, resetting the target position to the actual position and pressing move helps.
  •  
ELOG V3.1.4-7c3fd00